OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 15 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

OpenStudy (***[isuru]***):

hi, the first column t is an arithmetic sequence... 'cause the day s will varies as 1, 2, 3, 4, 5, ,6, .... so on until to 15 I think ya can understand that one... the second column is a geometric sequence.. 'cause if the weight is lose by 1/4 each day when t = 1 the weight is 32 g t = 2 the weight will be 32x(1/4) = 8 t = 3 the weight will be 8x(1/4) = 2 always the weight is multiplied by a common ratio to find the next weight.. so... that's a geometric progression Did ya get it ?

OpenStudy (anonymous):

ahhh yes I do

OpenStudy (***[isuru]***):

so... i guess ya have no problems this time :-)